Circular No. 8687

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006at AND 2006au Two supernovae have been discovered on unfiltered CCD survey im-

ages: 2006at by B. Dintinjana and H. Mikuž, Crni Vrh Observatory, on four frame taken with a 0.60-m f/3.3 Cichocki reflector in the course of their ‘Comet and Asteroid Search Program’ (PIKA); and 2006au by O. Trondal, P. Luckas, and M. Schwartz with the 0.35-m Tenagra telescope in Oslo (cf. IAUC 8674).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006at Mar. 8.055 13^{h}12^{m}41.^{s}11 +63^{o}16′45′′.4 17.1 8′′.0 E, 10′′.5 N
2006au Mar. 7.20 17 57 13.56 +12 11 03.2 17.2 17′′.0 W, 18′′.2 N

H. Yamaoka, Kyushu University; and H. Naito and N. Tokimasa, Nishi-Harima Astronomical Observatory, report that a rather noisy, lowresolution spectrum (range 420–690 nm; R = 1000 at 500 nm) taken of 2006at on Mar. 8.8 UT with the 2.0-m NAYUTA telescope shows a rather featureless blue (blackbody-like) continuum with possible very broad features (either absorption valleys or emission troughs, which suggests that it may be a very young supernova. Additional approximate magnitudes for 2006at: 1993 Mar. 19 UT, [19.6 (Digital Sky Survey, blue); 1997 Apr. 13, [20 (DSS, red); 2006 Feb. 12.074, [18.5 (PIKA R-band); Mar. 8.8, 16.7 (Yamaoka et al., slit-viewer unfiltered image). The host galaxy of 2006at is hard to identify on the DSS image because it is very diffuse; a SDSS color-composite image taken in 2001 also shows no point source down to mag ∼ 21. Additional approximate magnitudes for 2006au in UGC 11057: 2004 Sept. 17.15, [19.5; 2006 Mar. 10.65 UT, 17.4. The “Nearby Supernova Factory” collaboration reports that SN 2006au is a type-II supernova (details on CBET 427).

RS OPHIUCHI G. C. Anupama, Indian Institute of Astrophysics, Bangalore; and N. G.

Kantharia, National Center for Radio Astrophysics, Pune, report the lowfrequency radio detection of the current outburst of the recurrent nova RS Oph (cf. IAUC 8671) using the Giant Metrewave Radio Telescope (GMRT), at the following flux densities: Feb. 24.12 UT, 49.5 ± 0.5 mJy at 23.3 cm; Mar. 2.04, 55.4 ± 2.8 mJy at 28.3 cm, 50.0 ± 0.8 mJy at 23.3 cm, and 56.8 ± 0.3 mJy at 21.6 cm; Mar. 5.09, 48.4 ± 2.0 mJy at 49.2 cm. This is the first detection of the nova at wavelengths longer than 21 cm. Further GMRT observations at low frequencies are underway.

2006 March 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green